# Twinkle Twinkle Little Star, Roman Sees Where You Are: Predicting Exoplanet Transit Yields in the Rosette Nebula with the Nancy Grace Roman Space Telescope

### v1

@@ 5.2 Creating the Model Population
We generate a forward model of the stellar-planet population into our observed plane and then apply the survey parameters described in Section 2. For each star, we draw planetary systems from a spectral-type dependent occurrence rate model, map their age-agnostic radii to an age-dependent one, propagate line-of-sight dust to apparent magnitudes, and evaluate detectability with a stellar-variability noise model to determine transit search thresholds. We describe our procedure in further detail below.
5.2.1 Dust Extinction and Stellar Hosts
@@
We model line-of-sight dust using the dustmaps Python package (G. M. Green, 2018) with the Schlegel, Finkbeiner, and Davis (SFD) dustmap (1998, recalibrated by E. F. Schlafly & D. P. Finkbeiner 2011) sampled on a $0.5^{\circ}\times 0.5^{\circ}$ grid centered on the Rosette field. We convert E(BV)E(B-V) estimates to $A_{V}$ assuming the mean Galactic total-to-selective extinction ratio, $R_{V}=3.1$ (J. A. Cardelli et al., 1989). Our assumption is consistent with B. Fernandes et al. (2012), who found that most members in the core of the Rosette Nebula follow a normal extinction law, though localized deviations toward higher $R_{V}$ are possible in SFRs (adopting $R_{V}=4.0$ instead would only increase the mean $A_{\rm F146}$ by $\sim 0.4$ mag). We then convert $A_{V}$ the to F146 bandpass extinction using the E. L. Fitzpatrick (1999) extinction law evaluated at the filter’s effective wavelength. We find that the Rosette Nebula has a mean ${A}_{\mathrm{F146}}=1.25$ mag with a dispersion of 0.43 mag and a range of 1.784 mag. We randomly sample stars from Table 2 by spectral type, assigning extinction from a Gaussian centered at the field mean with $\sigma=0.43$.
5.2.2 Planetary Occurrence Rates
@@
Early planetary occurrence rate studies from the Kepler mission have shown that both the size and orbital period distribution of planets vary systematically with spectral type, with traditional core-accretion models also predicting the outward movement of the snow line with stellar mass and luminosity (G. M. Kennedy & S. J. Kenyon, 2008; G. Laughlin et al., 2004; C. Mordasini et al., 2012). For M dwarfs, planets are not only more numerous but occur at shorter orbital periods, often forming compact multi-planet systems with periods $\lessapprox 50$ days (C. D. Dressing & D. Charbonneau, 2015; G. D. Mulders et al., 2015; K. K. Hardegree-Ullman et al., 2019). In contrast, for planets orbiting FGK stars, the occurrence distribution shifts towards longer orbital periods, with a relative deficit of very short-period small planets and a rising frequency of giant planets at periods of hundreds of days (A. W. Howard et al., 2012; M. Kunimoto & S. Bryson, 2020).
@@
Recent modeling work by D. C. Hsu et al. (2019, 2020) account for these differences in planetary occurrence rates by using an Approximate Bayesian Framework (ABC) to infer occurrence distributions in the radius-period space from Kepler data. ABC is particularly well-suited for exoplanet demographics because it accommodates detection biases and incompleteness of transit surveys, while still allowing flexible parametric forms for the occurrence likelihood surface. In particular, D. C. Hsu et al. (2019) provided occurrence rate maps for FGK stars, whereas D. C. Hsu et al. (2020) extended this framework to M dwarfs, exploring both uniform and Dirichlet (a multivariate generalization of the $\beta$-distribution) priors. In this work, we adopt the FGK occurrence rates from these studies and employ the Dirichlet prior formulation for M dwarfs. Our choice of the Dirichlet prior is motivated by its ability to generate more realistic and physically consistent occurrence rate samples rather than a simple uniform prior (F. Komaki, 2012; A. Gelman et al., 2013).
@@
In bins with an asymmetric distribution where only the 83rd-percentile upper limits are reported, we instead construct a continuous occurrence surface using cubic radial basis function interpolation, a well-established method for smooth multi-dimensional approximation (e.g., M. J. D. Powell, 1992; M. D. Buhmann, 2003). Where the interpolated rate exceeds the reported upper limit, we adopt the latter, ensuring consistency with the statistical constraints while maintaining continuity across the radius-period space.
@@
Within this framework, each simulated star is assigned planetary companions drawn with empirically constrained occurrence distributions. Orbital inclinations (in particular, $\cos{i}$) for each system are then sampled from a uniform distribution (J. N. Winn, 2010), consistent with the co-planarity in observed orientations of multi-planet systems (J. Fang & J.-L. Margot, 2012; D. C. Fabrycky et al., 2014). Together, these steps produce a realistic synthetic planet population that underpins our injection-recovery analysis of the Rosette Nebula.
5.2.3 Radii Inflation
@@
Planetary radii are not static with age but contract as planets cool and lose residual heat from formation. Evolutionary models predict that young giant planets can have radii inflated by tens of percent compared to their older counterparts, with the degree of inflation depending on mass, composition, and formation pathways (I. Baraffe et al., 2003; J. J. Fortney et al., 2007; D. S. Spiegel & A. Burrows, 2012). Observational evidence reaffirms this idea: several young planetary systems host transiting planets with radii larger than expected for their masses and orbital separations (K2-33b and K2-25b, A. W. Mann et al. 2017; the multi-planet system V1298 Tau, T. J. David et al. 2019).
@@
To account for these effects in our model population, we adopt evolutionary cooling models from E. F. Linder et al. (2019), which provide radius-age scaling relations across a wide range of planetary masses. For each planet drawn from our occurrence rate distribution described in Section 5.2.2, we first determine age-agnostic mass-radius estimates from Forecaster (J. Chen & D. Kipping, 2017, 2016), which is a Bayesian forecasting model for objects covering nine orders-of-magnitude in mass. We then use these values to choose a mass-appropriate model and determine the scale factor given the planetary radius at 10 Myr. We present these scaling factors in Table 3.
@@
5.3 Generating Synthetic Lightcurves
@@
For each simulated host system, we generate transit models with the batman package (L. Kreidberg, 2015) at the survey cadence of 18 s, assuming uniform stellar disks and circular orbits. The injected transits are multiplied onto unity baselines and combined for multi-planet systems. We then add Gaussian noise consistent with the expected precision of each host star (Section 4.1), ensuring that the synthetic light curves reproduce survey noise characteristics while preserving transit depths. These lightcurves form the input for our injection-recovery pipeline.
5.4 Recovering Transiting Exoplanets
@@
To assess the detectability of injected planets, we process each synthetic lightcurve with the Box-Least-Squares (BLS) algorithm (G. Kovács et al., 2002; J. D. Hartman & G. A. Bakos, 2016), implemented within Lightkurve (Lightkurve Collaboration et al., 2018). For every host, we search across a grid of trial periods from 0.5-500 days and a range of transit durations appropriate for companions on both short and long period orbits. The maximum BLS periodogram peak corresponds to an initial candidate period, reference time of mid-transit, and detection signal-to-noise ratio for a given companion.
@@
Given that some simulated systems host multiple planets, we apply an iterative masking algorithm to disentangle overlapping signals: after each detection above our adopted $7\sigma$ threshold, the corresponding transit window is masked and BLS is re-run on the residuals. This procedure allows us to recover successive transiting planets in multi-planet systems rather than picking out the strongest signal. This iterative BLS framework is widely used in other transit surveys and algorithms (M. Hippke & R. Heller, 2019; J. Li et al., 2019; E. Tey et al., 2023), ensuring a well-validated recovery procedure. For every simulated planet, we record whether its orbit produced a transiting geometry and its recovered BLS statistic. We identify this as a detection if this statistic exceeds the $7\sigma$ threshold during one of the iterative passes. This yields a a consistent mapping between the injected synthetic population and the subset that would be observationally recoverable in this survey.
5.5 Results
@@
Figure 7: We present a 2D histogram with the distribution and density of transits recovered by host spectral type with one month of observations in radius-period space.
Figure 8: We show a simulated transiting population of exoplanets that Roman can detect in a month-long baseline. Red circles represent companions drawn from M dwarf occurrence rates whereas yellow circles represents companions drawn from FGK occurrence rates. We show the four innermost solar-system planets (dark blue circles) and the TRAPPIST-1 planets (purple circles). The sizes of each marker is scaled to the planet’s radius. We also show the habitability zone at 1 Gyr as estimated by R. K. Kopparapu et al. (2013) using MIST isochrones.
@@
To convert these detection fractions to yield counts, we scale by the underlying stellar population we infer from the IMF (Section 3.3). We estimate $409$ young FGK stars, $1290$ young M dwarfs, and an additional $1125$ BDs (Table 1). Applying the detection rates from Table 4, we predict that a one-month baseline would a yield $3\pm 1$ planets orbiting FGK hosts and $30^{+8}_{-7}$ planets orbiting M dwarfs. Likewise, we predict that a two-week baseline would yield $2\pm 1$ planets orbiting FGK hosts and $27\pm 7$ planets orbiting M dwarfs.
Despite being able to monitor BDs, given the absence of well-constrained BD planet detections, we do not extrapolate our adopted occurrence rates down to this mass regime. Given our survey’s capability of monitoring BDs at high S/N, detecting transiting companions around substellar hosts represents a novel and high-impact discovery space. Only a handful of BD-planet systems are known (e.g., G. Chauvin et al., 2004; D. P. Bennett et al., 2008; C. Han et al., 2013; Y. K. Jung et al., 2018), and their origins remain ambiguous between planetary and binary formation pathways (M. J. Payne & G. Lodato, 2007; B. P. Bowler, 2016). Given that young BDs are infrared bright, this survey of the Rosette Nebula offers excellent sensitivity to detect transiting planets orbiting these substellar hosts, potentially helping establish one of their first statistical samples in a young coeval population. Thus, detecting even a single transiting planet around a BD would represent a fundamentally new contribution to exoplanet demographics at the substellar boundary.
@@
Figure 8 illustrates one such simulated transiting planet population we recover in the Rosette Nebula, with the yield counts consistent with a month-long baseline, where the sizes of each marker are scaled to their radii. The distribution emphasizes the dominance of super-Earths and sub-Neptunes, but also highlights the presence of larger detectable planets at wider separations, underscoring the richness of this survey’s detectable population. We also overlay the 1 Gyr habitable zone-track from R. K. Kopparapu et al. (2013). We use Runaway Greenhouse and Maximum Greenhouse coefficients from their erratum-updated Table 3 combined with a 1 Gyr, solar-metallically MIST (A. Dotter, 2016; J. Choi et al., 2016; B. Paxton et al., 2013, 2015) isochrone track to determine this zone. Given that most transiting planets at this age are interior to the habitable zone, this underscores Roman’s potential to map out the evolutionary pathways of planets that may later enter habitable conditions. The Roman Space Telescope’s yields in the Rosette Nebula may therefore represent the progenitor population of future habitable-zone planets, allowing us to to directly probe the initial conditions and evolutionary pathways that support the formation of life.
6 Summary

### v2

@@ V.2 Creating the Model Population
We generate a forward model of the stellar-planet population into our observed plane and then apply the survey parameters described in Section II. For each star, we draw planetary systems from a spectral-type dependent occurrence rate model, map their age-agnostic radii to an age-dependent one, propagate line-of-sight dust to apparent magnitudes, and evaluate detectability with a stellar-variability noise model to determine transit search thresholds. We describe our procedure in further detail below.
V.2.1 Dust Extinction and Stellar Hosts
@@
We model line-of-sight dust using the dustmaps Python package (Green, 2018) with the Schlegel, Finkbeiner, and Davis (SFD) dustmap (1998, recalibrated by Schlafly and Finkbeiner 2011) sampled on a $0.5^{\circ}\times 0.5^{\circ}$ grid centered on the Rosette field. We convert E(BV)E(B-V) estimates to $A_{V}$ assuming the mean Galactic total-to-selective extinction ratio, $R_{V}=3.1$ (Cardelli et al., 1989). Our assumption is consistent with Fernandes et al. (2012), who found that most members in the core of the Rosette Nebula follow a normal extinction law, though localized deviations toward higher $R_{V}$ are possible in SFRs (adopting $R_{V}=4.0$ instead would only increase the mean $A_{\rm F146}$ by ${\sim}0.4$ mag). We then convert $A_{V}$ the to F146 bandpass extinction using the Fitzpatrick (1999) extinction law evaluated at the filter’s effective wavelength. We find that the Rosette Nebula has a mean ${A}_{\mathrm{F146}}=1.25$ mag with a dispersion of 0.43 mag and a range of 1.784 mag. We randomly sample stars from Table 2 by spectral type, assigning extinction from a Gaussian centered at the field mean with $\sigma=0.43$.
V.2.2 Planetary Occurrence Rates
@@
Early planetary occurrence rate studies from the Kepler mission have shown that both the size and orbital period distribution of planets vary systematically with spectral type, with traditional core-accretion models also predicting the outward movement of the snow line with stellar mass and luminosity (Kennedy and Kenyon, 2008; Laughlin et al., 2004; Mordasini et al., 2012). For M dwarfs, planets are not only more numerous but occur at shorter orbital periods, often forming compact multi-planet systems with periods $\lessapprox 50$ days (Dressing and Charbonneau, 2015; Mulders et al., 2015; Hardegree-Ullman et al., 2019). In contrast, for planets orbiting FGK stars, the occurrence distribution shifts towards longer orbital periods, with a relative deficit of very short-period small planets and a rising frequency of giant planets at periods of hundreds of days (Howard et al., 2012; Kunimoto and Bryson, 2020).
@@
Recent modeling work by Hsu et al. (2019, 2020) account for these differences in planetary occurrence rates by using an Approximate Bayesian Framework (ABC) to infer occurrence distributions in the radius-period space from Kepler data. ABC is particularly well-suited for exoplanet demographics because it accommodates detection biases and incompleteness of transit surveys, while still allowing flexible parametric forms for the occurrence likelihood surface. In particular, Hsu et al. (2019) provided occurrence rate maps for FGK stars, whereas Hsu et al. (2020) extended this framework to M dwarfs, exploring both uniform and Dirichlet (a multivariate generalization of the $\beta$-distribution) priors. In this work, we adopt the FGK occurrence rates from these studies and employ the Dirichlet prior formulation for M dwarfs. Our choice of the Dirichlet prior is motivated by its ability to generate more realistic and physically consistent occurrence rate samples rather than a simple uniform prior (Komaki, 2012; Gelman et al., 2013).
@@
In bins with an asymmetric distribution where only the 83rd-percentile upper limits are reported, we instead construct a continuous occurrence surface using cubic radial basis function interpolation, a well-established method for smooth multi-dimensional approximation (e.g., Powell, 1992; Buhmann, 2003). Where the interpolated rate exceeds the reported upper limit, we adopt the latter, ensuring consistency with the statistical constraints while maintaining continuity across the radius-period space.
@@
Within this framework, each simulated star is assigned planetary companions drawn with empirically constrained occurrence distributions. Orbital inclinations (in particular, $\cos{i}$) for each system are then sampled from a uniform distribution (Winn, 2010), consistent with the co-planarity in observed orientations of multi-planet systems (Fang and Margot, 2012; Fabrycky et al., 2014). Together, these steps produce a realistic synthetic planet population that underpins our injection-recovery analysis of the Rosette Nebula.
V.2.3 Radii Inflation
@@
Planetary radii are not static with age but contract as planets cool and lose residual heat from formation. Evolutionary models predict that young giant planets can have radii inflated by tens of percent compared to their older counterparts, with the degree of inflation depending on mass, composition, and formation pathways (Baraffe et al., 2003; Fortney et al., 2007; Spiegel and Burrows, 2012). Observational evidence reaffirms this idea: several young planetary systems host transiting planets with radii larger than expected for their masses and orbital separations (K2-33b and K2-25b, Mann et al. 2017; the multi-planet system V1298 Tau, David et al. 2019).
@@
To account for these effects in our model population, we adopt evolutionary cooling models from Linder et al. (2019), which provide radius-age scaling relations across a wide range of planetary masses. For each planet drawn from our occurrence rate distribution described in Section V.2.2, we first determine age-agnostic mass-radius estimates from Forecaster (Chen and Kipping, 2017, 2016), which is a Bayesian forecasting model for objects covering nine orders-of-magnitude in mass. We then use these values to choose a mass-appropriate model and determine the scale factor given the planetary radius at 10 Myr. We present these scaling factors in Table 3.
@@
V.3 Generating Synthetic Lightcurves
@@
For each simulated host system, we generate transit models with the batman package (Kreidberg, 2015) at the survey cadence of 18 s, assuming uniform stellar disks and circular orbits. The injected transits are multiplied onto unity baselines and combined for multi-planet systems. We then add Gaussian noise consistent with the expected precision of each host star (Section IV.1), ensuring that the synthetic light curves reproduce survey noise characteristics while preserving transit depths. These lightcurves form the input for our injection-recovery pipeline.
V.4 Recovering Transiting Exoplanets
@@
To assess the detectability of injected planets, we process each synthetic lightcurve with the Box-Least-Squares (BLS) algorithm (Kovács et al., 2002; Hartman and Bakos, 2016), implemented within Lightkurve (Lightkurve Collaboration et al., 2018). For every host, we search across a grid of trial periods from 0.5-500 days and a range of transit durations appropriate for companions on both short and long period orbits. The maximum BLS periodogram peak corresponds to an initial candidate period, reference time of mid-transit, and detection signal-to-noise ratio for a given companion.
@@
Given that some simulated systems host multiple planets, we apply an iterative masking algorithm to disentangle overlapping signals: after each detection above our adopted $7\sigma$ threshold, the corresponding transit window is masked and BLS is re-run on the residuals. This procedure allows us to recover successive transiting planets in multi-planet systems rather than picking out the strongest signal. This iterative BLS framework is widely used in other transit surveys and algorithms (Hippke and Heller, 2019; Li et al., 2019; Tey et al., 2023), ensuring a well-validated recovery procedure. For every simulated planet, we record whether its orbit produced a transiting geometry and its recovered BLS statistic. We identify this as a detection if this statistic exceeds the $7\sigma$ threshold during one of the iterative passes. This yields a a consistent mapping between the injected synthetic population and the subset that would be observationally recoverable in this survey.
[…]
Single transits have been proven viable for young planet discoveries and have traditionally served as a window into the population of long-period exoplanets (Wang et al., 2015; Foreman-Mackey et al., 2016). Recent confirmation of systems such as TOI-2180 b (Dalba et al., 2022) and TOI-2010 b (Mann et al., 2023) demonstrate that a single high-SNR event can provide crucial information via its transit duration and morphology to enable targeted follow-up (Namaste; Osborn et al., 2016). In the context of the Rosette Nebula, inclusion of single transits will be critical to capturing a larger diversity of the 10 Myr planetary census, especially for those orbital periods that might exceed our observation window.
Unlike TESS, which monitors nearby stars amenable to ground-based confirmation, the Rosette Nebula’s distance makes traditional follow up more challenging. JWST provides a compelling path for validating and characterizing high-priority candidates. The young age and inflated radii of these planets result in enhanced atmospheric scale heights favorable for transmission spectroscopy. Consequently, faint NIR targets in the Rosette Nebula are within JWST’s capabilities for both ephemeris refinement and atmospheric characterization using NIRSpec or NIRISS (Rustamkulov et al., 2023; Feinstein et al., 2023). Additionally, ground-based interferometry also offers a powerful validation pathway. While individual telescopes may lack Roman’s spatial resolution, interferometric facilities can achieve milliarcsecond-resolutions to identify unresolved stellar companions that could mimic transit signals.
V.5 Background Eclipsing Binaries
We assess the probability of background eclipsing binaries (BEBs) being misidentified as planetary transits. Roman’s WFI has a pixel scale of $0.11^{\prime\prime}$ (Akeson et al., 2019), comparable to that of HST. At the Rosette Nebula’s distance of 1400 pc, a conservative photometric aperture with radius $r_{\rm ap}=0.2^{\prime\prime}$ corresponds to a physical scale of ${\sim}280$ AU, enabling high spatial resolution. Given that the Rosette Nebula has a peak stellar density of $\Sigma=300$ stars pc-2, this also corresponds to an angular stellar density of $\rho_{*}=1.4\times 10^{-2}$ stars per arcsecond-2. Hence, the probability of a background star falling within our photometric aperture is,
$P_{\rm blend}$ $=\rho_{*}\times\pi r_{\rm ap}^{2}.$ (3)
This indicates that only ${\sim}0.08\%$ of our photometric apertures contain background stars. Of the small fraction of apertures containing background stars, the average occurrence rate of eclipsing binaries (EBs) across field populations is $f_{\rm EB}\approx 1.2\%$ (Prša et al., 2011). Furthermore, only shallow grazing eclipses can mimic planetary transits, further reducing $P_{\rm BEB}$. This exceptionally low BEB contamination rate is enabled by a combination Roman’s excellent spatial resolution and the Rosette Nebula’s modest stellar density compared to denser environments such as globular clusters (e.g., Gilliland et al., 2000).
However, we note that additional false positive mitigation will require standard vetting procedures. BEBs have been commonly identified by (1) centroid analysis during the transit, which can reveal photocentric shifts for blended sources (Batalha et al., 2013) and (2) secondary eclipse searches, as EBs produce secondary eclipses of comparable depth while planets produce much shallower (or undetectable) occultations. Our quoted yields therefore represent robust estimates of genuine planetary detections with minimal BEB contamination.
V.6 Results
@@
Figure 7: We present a 2D histogram with the distribution and density of transits recovered by host spectral type with one month of observations in radius-period space. We also show the marginal distributions of recovered planet radius for FGK stars (red) and M dwarfs (yellow).
Figure 8: We show a simulated transiting population of exoplanets that Roman can detect in a month-long baseline. Red circles represent companions drawn from M dwarf occurrence rates whereas yellow circles represents companions drawn from FGK occurrence rates. We show the four innermost solar-system planets (dark blue circles) and the TRAPPIST-1 planets (purple circles). The sizes of each marker is scaled to the planet’s radius. We also show the habitability zone at 1 Gyr as estimated by Kopparapu et al. (2013) using MIST isochrones.
@@
To convert these detection fractions to yield counts, we scale by the underlying stellar population we infer from the IMF (Section III.3). We estimate $409$ young FGK stars, $1290$ young M dwarfs, and an additional $1125$ BDs (Table 1). Applying the detection rates from Table 4, we predict that a one-month baseline would a yield $3\pm 1$ planets orbiting FGK hosts and $30^{+8}_{-7}$ planets orbiting M dwarfs. Likewise, we predict that a two-week baseline would yield $2\pm 1$ planets orbiting FGK hosts and $27\pm 7$ planets orbiting M dwarfs.
Despite being able to monitor BDs, given the absence of well-constrained BD planet detections, we do not extrapolate our adopted occurrence rates down to this mass regime. Given our survey’s capability of monitoring BDs at high S/N, detecting transiting companions around substellar hosts represents a novel and high-impact discovery space. Only a handful of BD-planet systems are known (e.g., Chauvin et al., 2004; Bennett et al., 2008; Han et al., 2013; Jung et al., 2018), and their origins remain ambiguous between planetary and binary formation pathways (Payne and Lodato, 2007; Bowler, 2016). Given that young BDs are infrared bright, this survey of the Rosette Nebula offers excellent sensitivity to detect transiting planets orbiting these substellar hosts, potentially helping establish one of their first statistical samples in a young coeval population. Thus, detecting even a single transiting planet around a BD would represent a fundamentally new contribution to exoplanet demographics at the substellar boundary.
@@
Our simulations also indicate that a majority of detections in a month-long baseline would multiple transits, with 64% for M dwarf planets and 54% for FGK star planets. These fractions reduce to 51% and 39% respectively for the two-week baseline, indicating that FGK planets benefit more from extended temporal coverage.
Figure 8 illustrates an example of a simulated transiting planet population we recover in the Rosette Nebula, with the yield counts consistent with a month-long baseline, where the sizes of each marker are scaled to their radii. The distribution emphasizes the dominance of super-Earths and sub-Neptunes, but also highlights the presence of larger detectable planets at wider separations, underscoring the richness of this survey’s detectable population. We also overlay the 1 Gyr habitable zone-track from Kopparapu et al. (2013). We use Runaway Greenhouse and Maximum Greenhouse coefficients from their erratum-updated Table 3 combined with a 1 Gyr, solar-metallically MIST (Dotter, 2016; Choi et al., 2016; Paxton et al., 2013, 2015) isochrone track to determine this zone. Given that most transiting planets at this age are interior to the habitable zone, this underscores Roman’s potential to map out the evolutionary pathways of planets that may later enter habitable conditions. The Roman Space Telescope’s yields in the Rosette Nebula may therefore represent the progenitor population of future habitable-zone planets, allowing us to to directly probe the initial conditions and evolutionary pathways that support the formation of life.
VI Summary